# Long-Wave Infrared Spintronic Poisson Bolometers with High Sensitivity

### v1

@@ Abstract
Abstract: High-sensitivity long-wave infrared (LWIR) detection is crucial for observing weak thermal radiation. Recently, the spintronic Poisson bolometer was proposed as a promising platform for uncooled infrared detection. The Poisson bolometer operates in a probabilistic regime dominated by Poissonian noise, establishing a new detection paradigm. In contrast to traditional analog detectors, where signal and noise are continuous currents or voltages, the Poisson bolometer has both signal and noise governed by Poissonian counting statistics regardless of the light source, with the mean count rate modulated by incident radiation. In this work, we integrate a broadband plasmonic absorber optimized for LWIR absorption onto a spintronic Poisson bolometer to enhance thermal coupling and temperature rise in the sensing layer. The plasmonic absorber achieves over 60% absorptance across the LWIR spectrum, matching the blackbody radiation peak at room temperature. The device exhibits a best noise-equivalent temperature difference (NEDT) of 35 mK at a 50 Hz frame rate and multiple results close to or below 100 mK, demonstrating room-temperature performance among the most sensitive uncooled LWIR detectors reported to date. This work advances uncooled infrared detection toward cryogenic-level sensitivity through the innovation of integrating spintronic materials and plasmonic materials, opening pathways to high-sensitivity LWIR sensing and imaging applications such as remote sensing, high-speed imaging, cryogenic system diagnostics, and industrial monitoring.
@@
Ziyi Yang, Sakshi Gupta, Jehan Shalabi, Daien He, Leif Bauer, Angshuman Deka, Zubin Jacob*
@@
Ziyi Yang, Jehan Shalabi, Dr. Daien He, Dr. Leif Bauer, Prof. Zubin Jacob
@@ 1 Introduction
High-sensitivity long-wave infrared (LWIR) detectors are essential for detecting weak thermal radiation. In many scenarios, the detectable infrared power is low due to radiative limits or measurement conditions. For example, low-temperature objects (<< 200 K) and low-emissivity surfaces emit faint infrared radiation [salihoglu2018graphene, xu2020high, HowDoesE48:online]. Moreover, targets with small temperature differences relative to their background require high thermal contrast resolution from detectors [jordan2024comparison, grimming2021ground]. Similarly, long-range imaging with high signal loss and high-speed imaging with short integration times both limit photon collection [flannigan2022mid, HighSpee43:online, pieper2017atmospheric]. These challenges underscore the need for highly sensitive LWIR detectors, which enable applications such as remote sensing of subtle geological variations, observation of cold celestial bodies, monitoring cryogenic systems, and imaging of low-emissivity or rapidly evolving scenes [rashman2020terrestrial, lawson2000brightness, wishnow1998long, landmann2019simultaneous, landmann2023high].
@@
State-of-the-art LWIR detectors can be classified into photon detectors and thermal detectors. Common LWIR photon detectors include type-II superlattice (T2SL) photodetectors typically based on InAs/GaSb or related antimonides [hu2025progress, zhou2023dual, FLIRA6780SLS, lee2021comparative, jiang2025high], quantum well infrared photodetectors (QWIP) employing GaAs/AlGaAs multiple quantum wells [lu2025lwir, ivanov2023qwip], mercury cadmium telluride (HgCdTe or MCT) photodiodes [zhang2023dark, varavin2020photodiodes], and superconducting nanowire single-photon detectors (SNSPDs) commonly fabricated from niobium nitride (NbN), niobium titanium nitride (NbTiN), or tungsten silicide (WSi) nanowires [chen2025sub, lita2022development, snspd_wsi]. Though demonstrating high response speed, high spectral selectivity, and sensitivity, these photon detectors are based on narrow-bandgap semiconductors or superconductors, fundamentally limited by thermally activated dark current, defect-mediated generation–recombination noise, or the temperature dependence of the superconducting gap [rogalski2003quantum, piotrowski2004uncooled, xue2023high, rogalski2023infrared, natarajan2012superconducting]. These noise mechanisms grow rapidly at room temperature and therefore necessitate cryogenic cooling to achieve high sensitivity.
@@
On the other hand, the most common LWIR thermal detectors are microbolometers, which measure radiation via temperature-induced change of a material’s electrical properties. Microbolometers typically employ vanadium oxide (VOx), amorphous silicon (a-Si), and are sometimes enhanced with metal-insulator-metal (MIM) absorbers to improve optical coupling [wu2023one, FLIRA655, fonseca2025new_amorSilicon, chen2020polarization]. Recent advances have also introduced thermomechanical detectors based on shape memory polymer (SMP) or perforated MIM absorber membrane as a promising new class of thermal detectors [adiyan2019shape, das2023thermodynamically]. However, both microbolometers and thermomechanical detectors face inherent constraints such as slow thermal time constants [yu2020low, rogalski2021trends], relatively large pixel footprints [yu2020low], and susceptibility to thermal or mechanical noise [vashaee2024thermoelectric, adiyan2019shape]. These limitations restrict response speed, resolution scaling, and ultimate sensitivity compared to cryogenically cooled photon detectors. As a result, achieving high sensitivity at room temperature remains an outstanding challenge for LWIR thermal detection. The most common figure of merit of the sensitivity of an infrared detector is the noise equivalent differential temperature (NEDT), or interchangeably noise equivalent temperature difference (NETD), representing the minimum temperature difference that can be sensed by the detector with a signal-to-noise ratio (SNR) equal to one [bianconi2020recent, rogalski2019infrared]. Visionary comparisons of the NEDT of state-of-the-art cooled and uncooled LWIR detectors are shown in Figure 1 (a),(b).
@@
Recently, the first Poisson bolometer operating at room temperature, called the spintronic ultrafast nanoscale (SUN) bolometer, was proposed [bauer2025exploiting]. The Poisson bolometer operates in a probabilistic regime dominated by Poissonian noise, establishing a novel detection paradigm. In contrast to traditional analog detectors, where signal and noise are continuous currents or voltages, the Poisson bolometer has both signal and noise governed by Poissonian counting statistics regardless of the light source, with the mean count rate modulated by incident radiation. This enables more effective separation of the noise and signal, potentially alleviating sensitivity limitations.
The Poisson bolometer in [bauer2025exploiting] was implemented using engineered spintronic materials, specifically a stochastic magnetic tunnel junction (MTJ). Thermally activated transitions between two discrete magnetization states were utilized to achieve a digital response to incident radiation. The reported spintronic Poisson bolometer demonstrates an NEDT of 103 mK at a 25 Hz frame rate, with an array of plasmonic nanoantennas incorporated onto the transduction layer to enhance the sensitivity. By comparison, a commercial VOx microbolometer-based camera (Teledyne FLIR A325sc) has an NEDT of 55 mK at 30 Hz. While achieving this sensitivity, the plasmonic absorber was designed to be highly absorptive to mid-wave infrared (MWIR) light, with only limited absorption in the LWIR, deviating from the radiation peak of a blackbody at room temperature.
In this work, we advance uncooled infrared detection toward cryogenic-level sensitivity by integrating a spintronic Poisson bolometer with a broadband LWIR plasmonic absorber to enhance thermal coupling. Through this innovation of integrating spintronic materials and plasmonic materials, we experimentally achieved a best NEDT of 35 mK at a 50Hz frame rate, and multiple NEDT values close to or below 100 mK, demonstrating room-temperature performance that ranks among the most sensitive uncooled LWIR detectors reported to date. The active pixel size of a spintronic Poisson bolometer is around 2 µm by 2 µm, potentially enabling high pixel density and resolution scaling. These results represent a significant improvement over previous spintronic Poisson bolometers and push the frontier of uncooled LWIR detection toward cryogenic-level sensitivity. We believe this work opens pathways to a broad range of high-sensitivity LWIR sensing and imaging applications such as remote sensing, high-speed imaging, cryogenic system diagnostics, and industrial monitoring.
@@ 2 Results
The spintronic Poisson bolometer we study here is composed of a broadband LWIR plasmonic absorber and a stochastic magnetic tunnel junction (MTJ). This integration enables a fundamentally new approach to detecting infrared radiation with tunable, magnetically controlled electrical readout, highlighting the innovation of spintronic materials in infrared detection. A simplified structure schematic is shown in Figure 1 (c). The magnetic tunnel junction comprises four primary layers: a magnetic sensing layer (i.e., free layer) based on CoFeB, a nonmagnetic insulating layer (i.e., tunnel barrier) based on MgO, a magnetic readout layer (i.e., fixed layer) also based on CoFeB, and a synthetic antiferromagnetic layer (SAF) consisting of Co/Pt multilayers separated by an Ir spacer. Detailed material stack information can be found in Supporting Information Figure S1. The MTJ exhibits tunnel magnetoresistance depending on the sensing layer’s magnetization orientation relative to that of the readout layer. The readout layer’s magnetization is fixed by the SAF layer located directly below, which is out-of-plane (perpendicular to the layers). The sensing layer has two stable magnetization states ($M_{1}$ and $M_{2}$) that are separated by an energy barrier. As a result, the device’s overall resistance depends on the sensing layer’s magnetic orientation. The resistance is low when the magnetizations of the sensing layer and the readout layer are parallel; the resistance is high when they are antiparallel. Experimentally, the resistance is read out by applying a bias voltage $V_{bias}$ to the spintronic Poisson bolometer and recording the bolometer’s voltage signal using an oscilloscope.
When no thermal radiation is incident, natural heat in the device causes discrete transitions between these two magnetization directions, which we refer to as dark counts. When thermal radiation is incident, the plasmonic absorption layer couples the heat to the sensing layer, which thermalizes through the MTJ. This increases the probability of transition in the MTJs sensing layer, leading to an increased count rate, as shown by Figure 1 (d). The relaxation time, i.e., interarrival time of the transitions, is governed by the Néel-Arrhenius law:
@@
where $\tau_{M_{1},(M_{2})}$ is the interarrival time, $\tau_{0}$ is the attempt time, $k_{B}$ is Boltzmann’s constant, and T is the temperature, and $E_{b}$ is the energy barrier:
@@
where $M$ is the magnetization, $H_{k}$ is the magnetic anisotropy field, V is the volume of the sensing layer, and H is the magnetic field present in the sensing layer. The offset in peak count rate from H = 0 is due to stray fields induced in the sensing layer by the nearby readout and SAF layer.
@@
Figure 2: Design and characterization of the LWIR plasmonic antenna array absorber. (a) Plasmonic absorber unit cell structure and dimensions. The absorber is a metal-insulator-metal (MIM) type of metasurface with circular Ti nanoantennas on Ge-Ti layers. The dimensions are engineered to achieve enhanced broadband LWIR absorption. (b) SEM image of the fabricated spintronic Poisson bolometer device. The MTJ nanopillar bridges two top gold contacts and two bottom gold contacts. The top contact strip area above the MTJ is covered with the plasmonic absorber. The left inset shows the plasmonic nanoantenna array. The MTJ is located near the nanoantenna to obtain higher heat transfer. (c) Left vertical axis: Simulated and measured absorptance of the plasmonic absorber. The design enables broadband LWIR absorption higher than 60% with a peak around 11 µm wavelength. Right vertical axis: Simulated blackbody spectral radiance at 300 K. At room temperature, the blackbody radiation peaks at the LWIR spectrum (red shaded region). The absorption of the plasmonic antenna array peaks at the same region to maximize the temperature increase in the sensing layer of the MTJ.
@@
To optimize thermal coupling and enhance the temperature rise in the device’s sensing layer, we designed and integrated a broadband LWIR plasmonic absorber onto a spintronic Poisson bolometer. The plasmonic absorber adopted here is a metal-insulator-metal (MIM) type of metasurface designed based on the standing wave model [nath2015far, to2020detailed, zhou2021ultra, liu2010infrared]. The plasmonic absorber unit cell structure and dimensions are shown in Figure 2 (a). The absorber is composed of periodic disk-shaped titanium nanoantennas and a bottom titanium layer, separated by a germanium dielectric layer. Titanium is chosen as both the antenna and the bottom layer material due to its lossy and highly reflective properties at LWIR to achieve broadband absorption [ding2016broadband]. Germanium is chosen as the dielectric layer due to its high refractive index ($n=4$). The antenna diameter, unit cell dimension, and layer thicknesses are engineered to achieve enhanced broadband LWIR absorption. Figure 2 (b) shows an SEM image of the fabricated spintronic Poisson bolometer device. The MTJ nanopillar bridges two top gold contacts and two bottom gold contacts. The top contact strip area above the MTJ is covered with the plasmonic absorber. The left inset shows the plasmonic nanoantenna array. The MTJ is located near the nanoantenna to obtain higher heat transfer.
@@
Figure 2 (c) shows the simulated and measured absorptance of the plasmonic absorber. This design enables broadband LWIR absorption higher than 60% across the entire LWIR spectrum and peaks around 11 µm wavelength. The simulated blackbody spectral radiance at 300 K suggests that the blackbody radiation peaks at the LWIR spectrum (red shaded region). The absorption of the plasmonic antenna array peaks at the same region to maximize the temperature increase in the sensing layer of the MTJ. Compared to the broadband MWIR plasmonic absorber design in [bauer2025exploiting], the broadband LWIR plasmonic absorber in this paper enables around 4 times the original temperature increase in the sensing layer according to the heat transfer simulation by COMSOL (see Supporting Information Table 1).
@@
We first set the blackbody temperature to 293.15 K, the same as the environmental temperature, and characterize the baseline transitions. Figure 3 (c) shows the voltage waveforms, where clear transitions between two voltage levels can be seen and a continuous tuning of the count rate by the applied field is observed. We take a histogram of the waveform and set voltage thresholds accordingly to calculate the number of counts or count rates. Figure 3 (d) shows the field dependence of the baseline count rate with 20 ms integration time. The dashed line suggests a good fit to the Néel-Arrhenius.
Figure 4: NETDT measurement results. (a) Measured NEDT of a FLIR A325sc LWIR camera. This result is measured with an f/0.6 camera lens and 30 Hz frame rate. No external LWIR lens is used for the camera. The distance between the camera lens and the blackbody is 5 cm. The 79 mK result is close to the datasheet’s 50 mK sensitivity specification, validating the measurement setup’s effectiveness, and is used as a benchmark of a commercial uncooled LWIR detector based on VOx microbolometers. The error bars in this plot and the following plots represent the standard deviation of 5 frames. (b) NEDT result measured with dT = 0.3 K. For this applied field and bias voltage, the device shows a linear response between 24.7 and 25.3 °C. Afterward, the count rate stops increasing due to device saturation. (c) Repeated NEDT measurement with dT = 0.1 K with the same applied field and bias voltage as (b), but with a focus on the linear response region to achieve the best possible sensitivity. The result shows an improved NEDT of 35 mK. (d) Repeated NEDT measurement with decreased bias voltage and slightly increased applied field. The result shows a decreased count rate and higher NEDT due to both the current-induced spin-transfer torque effect and the field dependence of the count rate. With the bias voltage and the applied field properly selected, the spintronic Poisson bolometer is able to achieve close to, well below 100 mK NEDT in multiple measurements on different days. (e)-(f) NEDT measured with negative bias voltage. The result shows the tunability of device performance with the bias voltage and applied field. Each of these 6 measurements is done on a different day. The integration time for (b) - (d) is 20 ms (50 Hz frame rate). The integration time for (e) and (f) is 40 ms (25 Hz frame rate), all comparable to the frame rate of state-of-the-art commercial LWIR cameras.
To validate the measurement setup’s effectiveness and benchmark a commercial uncooled LWIR detector, we measure the NEDT of a VOx microbolometer-based camera (Teledyne FLIR A325sc). The NEDT is defined as:
@@
where $C_{n}$ is the count rate noise, i.e., standard deviation of detector signal for a blackbody at the temperature of interest, and $dC/dT$ is the responsivity, i.e., count rate change per kelvin change in blackbody temperature. For all measurements in this paper, the standard deviation is taken from 5 frames. Figure 4 (a) shows the NEDT result of the FLIR A325sc LWIR camera. The 79 mK result is close to the datasheet’s 50 mK sensitivity specification. This result is measured with an f/0.6 camera lens and 30 Hz frame rate. No external LWIR lens is used for the camera. The distance between the camera lens and the blackbody is 5 cm. Following this, we did five NEDT measurements of a spintronic Poisson bolometer on five different days. The results are shown in Figure 4 (b)-(f).

### v3

# Enhanced Sensitivity to Blackbody Radiation in Spintronic Poisson Bolometers

@@ Abstract
Abstract: High-sensitivity long-wave infrared (LWIR) detection is crucial for observing weak thermal radiation. Recently, the Poisson bolometer is proposed as a fundamentally new platform for uncooled infrared detection. In contrast to traditional analog detectors, where signal and noise are continuous currents or voltages, the Poisson bolometer has both signal and noise governed by Poissonian counting statistics regardless of the light source. In this work, uncooled infrared detection is advanced toward cryogenic-level sensitivity through the integration of spintronic materials and plasmonic materials. Specifically, a plasmonic absorber optimized for broadband LWIR absorption is experimentally integrated with a spintronic Poisson bolometer to enhance thermal coupling and temperature rise in the sensing layer. The plasmonic absorber exhibits an absorptance exceeding 60% across the LWIR spectrum, matching the peak of room-temperature blackbody radiation. The device achieves a noise-equivalent temperature difference (NEDT) of 35 mK at a 50 Hz frame rate and multiple results close to or below 100 mK, demonstrating room-temperature performance comparable to the most sensitive uncooled LWIR detectors reported to date. This work opens up pathways to high-sensitivity LWIR detection and imaging applications, including remote sensing, high-speed imaging, and industrial monitoring, while eliminating bulky and expensive cooling requirements.
@@
Ziyi Yang, Sakshi Gupta, Jehan Shalabi, Daien He, Leif Bauer, Mohamed A. Mousa, Angshuman Deka, Zubin Jacob*
@@
Ziyi Yang, Jehan Shalabi, Dr. Daien He, Dr. Leif Bauer, Mohamed A. Mousa, Prof. Zubin Jacob
@@ 1 Introduction
High-sensitivity long-wave infrared (LWIR) detectors are essential for detecting weak thermal radiation [43]. In many scenarios, the detectable infrared power is low due to radiative limits or measurement conditions. For example, low-temperature objects (<< 200 K) and low-emissivity surfaces emit faint infrared radiation [44, 51, 20]. Moreover, targets with small temperature differences relative to their background require high thermal contrast resolution from detectors [24, 17]. Similarly, long-range imaging with high signal loss and high-speed imaging with short integration times both limit photon collection [12, 19]. These challenges underscore the need for highly sensitive LWIR detectors, which enable applications such as remote sensing of subtle geological variations, observation of cold celestial bodies, monitoring cryogenic systems, and imaging of low-emissivity or rapidly evolving scenes [39, 27, 49, 25, 26].
@@
State-of-the-art LWIR detectors can be classified into photon detectors and thermal detectors. Common LWIR photon detectors include type-II superlattice (T2SL) photodetectors typically based on InAs/GaSb or related antimonides [21, 56, 31, 28, 23], quantum well infrared photodetectors (QWIP) employing GaAs/AlGaAs multiple quantum wells [32, 22], mercury cadmium telluride (HgCdTe or MCT) photodiodes [55, 46], and superconducting nanowire single-photon detectors (SNSPDs) commonly fabricated from niobium nitride (NbN), niobium titanium nitride (NbTiN), or tungsten silicide (WSi) nanowires [8, 29, 48]. Though demonstrating high response speed, high spectral selectivity, and sensitivity, these photon detectors are based on narrow-bandgap semiconductors or superconductors, fundamentally limited by thermally activated dark current, defect-mediated generation–recombination noise, or the temperature dependence of the superconducting gap [42, 38, 52, 40, 36]. These noise mechanisms grow rapidly at room temperature and therefore necessitate cryogenic cooling to achieve high sensitivity.
@@
On the other hand, the most common LWIR thermal detectors are microbolometers, which measure radiation via temperature-induced change of a material’s electrical properties [43]. Microbolometers typically employ vanadium oxide (VOx), amorphous silicon (a-Si), and are sometimes enhanced with metal-insulator-metal (MIM) absorbers to improve optical coupling [50, 13, 14, 7]. Recent advances have also introduced thermomechanical detectors based on shape memory polymer (SMP) or perforated MIM absorber membrane as a promising new class of thermal detectors [2, 10]. However, both microbolometers and thermomechanical detectors face inherent constraints such as slow thermal time constants [54, 41], relatively large pixel footprints [54], and susceptibility to thermal or mechanical noise [47, 2]. These limitations restrict response speed, resolution scaling, and ultimate sensitivity compared to cryogenically cooled photon detectors. As a result, achieving high sensitivity at room temperature remains an outstanding challenge for LWIR thermal detection. The most common figure of merit of the sensitivity of an infrared detector is the noise equivalent differential temperature (NEDT), or interchangeably noise equivalent temperature difference (NETD), representing the minimum temperature difference that can be sensed by the detector with a signal-to-noise ratio (SNR) equal to one [4, 43]. Visionary comparisons of the NEDT of state-of-the-art cooled and uncooled LWIR detectors are shown in Figure 1 (a),(b).
@@
Recently, the first Poisson bolometer operating at room temperature, called the spintronic ultrafast nanoscale bolometer, was proposed [3, 35]. The Poisson bolometer operates in a probabilistic regime dominated by Poissonian noise, establishing a novel detection paradigm [53]. In contrast to traditional analog detectors, where signal and noise are continuous currents or voltages, the Poisson bolometer has both signal and noise governed by Poissonian counting statistics regardless of the light source, with the mean count rate modulated by incident radiation. Poisson‑counting detection has advantages in low‑signal regimes relative to analog detection with Gaussian noise, as shown in photon counting literature where event‑based detection reduces sensitivity to continuous noise and enhances low‑light sensitivity [15].
The Poisson bolometer in [3] was implemented using engineered spintronic materials, specifically a stochastic magnetic tunnel junction (MTJ) based on CoFeB/MgO/CoFeB and a synthetic antiferromagnetic layer (SAF). Thermally activated transitions between two discrete magnetization states were utilized to achieve a digital response to incident radiation. The reported spintronic Poisson bolometer demonstrates an NEDT of 103 mK at a 25 Hz frame rate, with an array of plasmonic nanoantennas incorporated onto the transduction layer to enhance sensitivity. While achieving this sensitivity, the plasmonic absorber was designed to be highly absorptive to mid-wave infrared (MWIR) light, with only limited absorption in the LWIR, deviating from the radiation peak of a blackbody at room temperature.
In this work, we advance uncooled infrared detection toward cryogenic-level sensitivity by integrating a spintronic Poisson bolometer (SPB) with a broadband LWIR plasmonic absorber to enhance thermal coupling. Through this innovation of integrating spintronic materials and plasmonic materials, we experimentally achieved a best NEDT of 35 mK at a 50 Hz frame rate, and multiple NEDT values close to or below 100 mK, demonstrating room-temperature performance that ranks among the most sensitive uncooled LWIR detectors reported to date. The active pixel size of a spintronic Poisson bolometer is approximately 2 µm by 2 µm, potentially allowing for high pixel density and resolution scaling. These results represent a significant improvement over previous spintronic Poisson bolometers and push the frontier of uncooled LWIR detection toward cryogenic-level sensitivity. We believe that this work opens the way to a broad range of high-sensitivity LWIR sensing and imaging applications, including remote sensing, high-speed imaging, cryogenic system diagnostics, and industrial monitoring.
@@ 2 Results
The spintronic Poisson bolometer we study here is composed of a broadband LWIR plasmonic absorber and a stochastic magnetic tunnel junction (MTJ). This integration enables a fundamentally new approach to detecting thermal radiation with tunable, magnetically controlled electrical readout, highlighting the innovation of spintronic materials in infrared detection. A simplified structure schematic is shown in Figure 1 (c). The magnetic tunnel junction comprises four primary layers: a magnetic sensing layer (i.e., free layer) based on CoFeB, a nonmagnetic insulating layer (i.e., tunnel barrier) based on MgO, a magnetic readout layer (i.e., fixed layer) also based on CoFeB, and a synthetic antiferromagnetic layer (SAF) consisting of Co/Pt multilayers separated by an Ir spacer. Detailed material stack information can be found in Supporting Information Figure S1. The MTJ exhibits tunnel magnetoresistance depending on the sensing layer’s magnetization orientation relative to that of the readout layer. The readout layer’s magnetization is fixed by the SAF layer located directly below, which is out-of-plane (perpendicular to the layers). The sensing layer has two stable magnetization states ($M_{1}$ and $M_{2}$) that are separated by an energy barrier. As a result, the device’s overall resistance depends on the sensing layer’s magnetic orientation. The resistance is low when the magnetizations of the sensing layer and the readout layer are parallel; the resistance is high when they are antiparallel. Experimentally, the resistance is read out by applying a bias voltage $V_{bias}$ to the spintronic Poisson bolometer and recording the bolometer’s voltage signal using an oscilloscope.
When no thermal radiation is incident, natural heat in the device causes discrete transitions between these two magnetization directions, which we refer to as dark counts. When thermal radiation is incident, the plasmonic absorption layer couples the heat to the sensing layer, which thermalizes through the MTJ. This increases the probability of transition in the MTJs sensing layer, leading to an increased count rate, as shown by Figure 1 (d). The mean of the relaxation times, i.e., interarrival times, of the stochastic transitions $\tau_{M_{1},(M_{2})}$, is governed by the Néel-Arrhenius law [9]:
@@
where $\tau_{0}$ is the attempt time, $k_{B}$ is Boltzmann’s constant, and $T$ is the temperature of the sensing layer, and $E_{b}$ is the energy barrier:
@@
where $M$ is the magnetization, $V$ is the volume of the sensing layer, $H_{k}$ is the magnetic anisotropy field, and $H$ is the magnetic field present in the sensing layer. The Néel–Arrhenius law can alternatively be written in terms of the rate of transitions, i.e., count rate:
$\lambda(T)=\lambda_{0}\exp\Bigg(-\frac{E_{b}}{k_{B}T}\Bigg)$ (3)
where $\lambda_{0}=1/\tau_{0}$ is the attempt rate. On the other hand, the relaxation time is known to follow an exponential distribution [18, 1], and the counts follow a Poisson distribution [16, 5]. Therefore, during an observation window, or integration time $\Delta t$, the number of counts $N(\Delta t)$ is:
$N(\Delta t)=\mathrm{Poisson}(\mu)$ (4)
The mean of the Poisson distribution is $\mu=\lambda\Delta t$. When no light is incident, assume the temperature of the free layer is $T_{0}$, then the count rate is:
$N_{0}(\Delta t)=\mathrm{Poisson}\Big(\mu_{0}=\lambda_{0}\Delta t\,e^{-E_{b}/(k_{B}T_{0})}\Big)$ (5)
According to [3], the count rate increases linearly with incident power; therefore, the temperature increase $\Delta T=T-T_{0}$ can be assumed to be linear with the incident power $P_{\rm in}$ when $\Delta T\ll T_{0}$:
$T=T_{0}+\eta P_{\rm in}$ (6)
where $\eta$ is the thermal responsivity. The count rate becomes:
$N_{1}(\Delta t)=\mathrm{Poisson}\Big(\mu_{1}=\lambda_{0}\Delta t\,e^{-E_{b}/(k_{B}(T_{0}+\eta P_{\rm in}))}\Big)$ (7)
The mean count rate shifts from $\mu_{0}$ to $\mu_{1}$ due to the incident light, and this is the key to estimating $P_{\rm in}$. For a Poisson process, the maximum likelihood estimator (MLE) of the count rate from the number of counts observed in $\Delta t$ is
$\hat{\lambda}=\frac{N(\Delta t)}{\Delta t}$ (8)
The estimator of $P_{\rm in}$ can be obtained by inverting the count rate equation (7):
$\hat{P}_{\rm in}=\frac{1}{\eta}\left[\frac{E_{b}}{k_{B}}\left(\ln\lambda_{0}-\ln\hat{\lambda}\right)^{-1}-T_{0}\right]$ (9)
@@
Figure 2: Design and characterization of the LWIR plasmonic antenna array absorber. (a) Plasmonic absorber unit cell structure and dimensions. The absorber is a metal-insulator-metal (MIM) type of metasurface with circular Ti nanoantennas on Ge-Ti layers. The dimensions are engineered to achieve enhanced broadband LWIR absorption. (b) SEM image of the fabricated spintronic Poisson bolometer. The MTJ nanopillar bridges two top gold contacts and two bottom gold contacts. The top contact strip area above the MTJ is covered with the plasmonic absorber. The left inset shows the plasmonic nanoantenna array. The MTJ is located near the nanoantenna to obtain higher heat transfer. (c) Left vertical axis: Simulated and measured absorptance of the plasmonic absorber. The design enables broadband LWIR absorption higher than 60% with a peak around 11 µm wavelength. Right vertical axis: Simulated blackbody spectral radiance at 300 K. At room temperature, the blackbody radiation peaks at the 7-14 µm LWIR spectrum (red shaded region). The absorption of the plasmonic antenna array peaks at the same region to maximize the temperature increase in the sensing layer of the MTJ.
@@
To optimize thermal coupling and enhance the temperature rise in the device’s sensing layer, we integrated a broadband LWIR plasmonic absorber onto a spintronic Poisson bolometer. The plasmonic absorber adopted here is a metal-insulator-metal (MIM) type of metasurface designed based on the standing wave model [37, 45, 57, 30]. The plasmonic absorber unit cell structure and dimensions are shown in Figure 2 (a). The absorber is composed of periodic disk-shaped titanium nanoantennas and a bottom titanium layer, separated by a germanium dielectric layer. Titanium is chosen as both the antenna and the bottom layer material due to its lossy and highly reflective properties at LWIR to achieve broadband absorption [11]. Germanium is chosen as the dielectric layer due to its high refractive index ($n=4$). The antenna diameter, unit cell dimension, and layer thicknesses are engineered to achieve enhanced broadband LWIR absorption. Figure 2 (b) shows an SEM image of the fabricated spintronic Poisson bolometer device. The MTJ nanopillar bridges two top gold contacts and two bottom gold contacts. The top contact strip area above the MTJ is covered with the plasmonic absorber. The left inset shows the plasmonic nanoantenna array. The MTJ is located near the nanoantenna to obtain higher heat transfer.
@@
Figure 2 (c) shows the simulated and measured absorptance of the plasmonic absorber. This design enables broadband LWIR absorption higher than 60% across the entire LWIR spectrum and peaks around 11 µm wavelength. The simulated blackbody spectral radiance at 300 K suggests that the blackbody radiation peaks at the 7-14 µm LWIR spectrum (red shaded region). The absorption of the plasmonic antenna array peaks at the same region to maximize the temperature increase in the sensing layer of the MTJ. Compared to the broadband MWIR plasmonic absorber in [3], the broadband LWIR plasmonic absorber in this paper enables around 4 times the original temperature increase in the sensing layer according to the heat transfer simulation by COMSOL (see Supporting Information Table 1).
@@
We first set the blackbody temperature to 293.15 K, the same as the environmental temperature, and characterize the baseline transitions, i.e., stabilized dark counts pinned by the blackbody at room temperature. Figure 3 (c) shows the voltage waveforms, where clear transitions between two voltage levels can be seen, and a continuous tuning of the count rate by the applied field is observed. We take a histogram of the waveform and set voltage thresholds accordingly to calculate the number of counts or count rates. Figure 3 (d) shows the field dependence of the baseline count rate with 20 ms integration time. The dashed line suggests a good fit to the Néel-Arrhenius with an offset in the peak count rate from $H=0$ due to stray fields induced in the sensing layer by the nearby readout and SAF layer. Since the spintronic Poisson bolometer operates via discrete transition events, its dominant noise source is dark counts. The dark count rate plays a critical role in determining sensitivity, and its controllability through the applied magnetic field offers a direct mechanism for tuning and engineering device performance.
Figure 4: NEDT measurement results. (a) Measured NEDT of a VOx microbolometer-based LWIR camera (FLIR A325sc). This result is measured with an f/0.6 camera lens and 30 Hz frame rate. No external LWIR lens is used for the camera. The distance between the camera lens and the blackbody is 5 cm. The 79 mK result is close to the datasheet’s 50 mK sensitivity specification, validating the measurement setup’s effectiveness, and is used as a benchmark of a commercial uncooled LWIR detector based on VOx microbolometers. The error bars in this plot and the following plots represent the standard deviation of 5 frames. (b) NEDT result measured with dT = 0.3 K. For this applied field and bias voltage, the device shows a linear response between 24.7 and 25.3 °C. Afterward, the count rate stops increasing due to device saturation. (c) Repeated NEDT measurement with dT = 0.1 K with the same applied field and bias voltage as (b), but with a focus on the linear response region to achieve the best possible sensitivity. The result shows an improved NEDT of 35 mK. (d) Repeated NEDT measurement with decreased bias voltage and slightly increased applied field. The result shows a decreased count rate and higher NEDT due to both the current-induced spin-transfer torque effect and the field dependence of the count rate. With the bias voltage and the applied field properly selected, the spintronic Poisson bolometer can achieve close to, well below 100 mK NEDT in multiple measurements on different days. (e)-(f) NEDT measured with negative bias voltage. The result shows the tunability of device performance with the bias voltage and applied field. Each of these 6 measurements is done on a different day. The integration time for (b) - (d) is 20 ms (50 Hz frame rate). The integration time for (e) and (f) is 40 ms (25 Hz frame rate), all comparable to the frame rate of state-of-the-art commercial LWIR cameras.
To validate the measurement setup’s effectiveness and benchmark a representative uncooled LWIR detector, we measure the NEDT of a VOx microbolometer-based camera (Teledyne FLIR A325sc). The NEDT is defined as:
@@
where $C_{n}$ is the count rate noise, i.e., standard deviation of detector signal for a blackbody at the temperature of interest, and $dC/dT$ is the responsivity, i.e., count rate change per kelvin change in blackbody temperature. For all measurements in this paper, the standard deviation is taken from 5 frames. Figure 4 (a) shows the NEDT result of the VOx microbolometer-based LWIR camera. The 79 mK result is close to the datasheet’s 50 mK sensitivity specification. This result is measured with an f/0.6 camera lens and 30 Hz frame rate. No external LWIR lens is used for the camera. The distance between the camera lens and the blackbody is 5 cm. Following this, we did five NEDT measurements of a spintronic Poisson bolometer on five different days. The results are shown in Figure 4 (b)-(f).